A 33 g ice cube at −14◦C is dropped into a
container of water at 0◦C.
How much water freezes onto the ice? The
specific heat of ice is 0.5 cal/g ·
◦ C and its heat
of fusion of is 80 cal/g.
Answer in units of g.
1
Expert's answer
2020-04-05T14:51:12-0400

The equation of heat balance says

"c_{ice}m_{ice}\Delta T=\lambda m_{water}."

Hence, the amount of water that freezes

"m_{water}=\frac{c_{ice}m_{ice}\Delta T}{\lambda}""=\frac{0.5\times 33\times 14}{80}=2.9\:\rm g"